# FedAdamW: A Communication-Efficient Optimizer with Convergence and Generalization Guarantees for Federated Large Models

### v1

@@ Introduction
With the rapid growth of data and rising concerns over user privacy, traditional centralized training paradigms have become inadequate. Federated Learning (FL) (McMahan et al. 2017) offers a scalable and privacy-preserving framework that enables collaborative model training across decentralized clients without sharing raw data (Liu et al. 2024, 2025b, 2025a). As data becomes increasingly siloed, FL is a practical solution for large-scale distributed deep learning.
However, recent trends in model design-particularly the rise of large-scale architectures such as GPT (Radford et al. 2018), RoBERTa (Liu et al. 2019), and Vision Transformers (ViT) (Dosovitskiy et al. 2020)—pose new challenges for existing FL algorithms. Specifically, the widely-used FedAvg algorithm, which relies on stochastic gradient descent (SGD(Bottou 2010) in local, struggles to efficiently train Transformer models. This is due to the slow convergence and poor adaptivity of SGD in Transformer models (Zhang et al. 2024a; Liu et al. 2025a), which have more complex architectures compared to CNNs. For example, components such as query, key, and value often require different learning rates to be trained effectively (Zhang et al. 2024a). In contrast, AdamW (Loshchilov, Hutter et al. 2017), an adaptive optimizer with decoupled weight decay, has demonstrated superior performance in centralized training of large models based on Transformer (Vaswani et al. 2017; Liu et al. 2019), offering faster convergence and improved generalization, compared to Adam (Kingma and Ba 2014) and SGD.
Empirically, we also observe this advantage in FL: as shown in Figure 1, local training with AdamW (Local AdamW) converges significantly faster than Local SGD  (McMahan et al. 2017) for training various Transformer models . However, naively applying AdamW in FL leads to the following new challenges:
@@
Figure 1: Performance of Local SGD and Local AdamW. For training ViT-Base, GPT2, and BERT (Liu et al. 2019), we carefully tune the learning rate. For training all these Transformer models, Local SGD is still significantly worse than Local AdamW.
@@
  • We propose FedAdamW, a principled FL algorithm tailored for adaptive optimizers. To address the above challenges, FedAdamW integrates global update estimate into local updates to mitigate overfitting and improve consistency. Inspired by the Hessian structure, we design a communication-efficient aggregation strategy that communicates the mean of second-moment estimates across clients.
@@
$\bullet$ Heterogeneity Issues in Federated Learning. Data heterogeneity across clients is a fundamental challenge in FL. A range of algorithms have been proposed to mitigate the adverse effects of non-i.i.d. data distributions. For example, FedProx (Li et al. 2020) introduces a proximal term to restrict local updates; SCAFFOLD (Karimireddy et al. 2020) applies control variates to correct client drift; and FedCM (Xu et al. 2021) leverages client momentum to stabilize updates. These methods are predominantly built upon the SGD optimizer, and thus inherit its limitations when optimizing large-scale or deep Transformer-based models.
@@
$\bullet$ Adaptive Optimization in Centralized Settings. Adaptive gradient methods have demonstrated superior empirical performance over SGD in centralized settings, particularly for deep neural networks. Pioneering works include Adagrad (Duchi, Hazan, and Singer 2011), Adadelta (Zeiler 2012), Adam (Kingma and Ba 2014), AMSGrad (Reddi, Kale, and Kumar 2019), and AdamW (Loshchilov, Hutter et al. 2017). AdamW, in particular, decouples weight decay from gradient updates, offering improved generalization and training stability—attributes especially critical for Transformer models (Liu et al. 2019; Zhang et al. 2024a).
@@
$\bullet$ Adaptive Optimization in Federated Learning. Recent efforts have explored integrating adaptive methods into FL. FedOpt (Reddi et al. 2020) incorporates server-side adaptivity using Adam and Yogi. FAFED (Wu et al. 2023) aggregates both the first- and second-moment estimates of Adam across clients to stabilize training. FedAMS  (Chen, Li, and Li 2020) shows that averaging the second-moment estimate of Adam is crucial to prevent divergence. More recently, Sun et al. (2023) proposed to only aggregate the second-moment estimate to reduce communication overhead. However, these works only conducted experiments on CNN models. These studies are all based on Adam, which performs poorly with large weight decay.
$\bullet$ Our Contribution for Large Models. We propose FedAdamW, the first federated learning algorithm based on AdamW tailored for Transformer models. Our method integrates a global-local update alignment strategy and a communication-efficient aggregation of second-moment estimates. Furthermore, we provide theoretical convergence and PAC-Bayesian generalization analysis, offering new insights into adaptive optimization in FL.
@@
Challenges of AdamW in FL
Despite the widespread use of AdamW (Loshchilov, Hutter et al. 2017; Vaswani et al. 2017) in centralized deep learning, its adaptation to federated settings remains largely unexplored. In this section, we analyze three fundamental challenges that hinder its effectiveness in FL settings.
where $\boldsymbol{v}^{r,k}_{i}$ denotes the second-moment estimate maintained by client $i$ at local step $k$ of round $r$, $\boldsymbol{g}^{r,k}_{i}$ is the stochastic gradient, $\beta_{2}=0.999$ is the exponential decay rate for the second moment, and $\odot$ represents the element-wise (Hadamard) product in Algorithm 1. In FL, data heterogeneity leads to gradient heterogeneity. The squared stochastic gradients $\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$ in Eq. (2) amplify the variance of $\boldsymbol{v}$ across clients in Figure 2 (a). This can cause instability and inefficient aggregation, especially when using non-i.i.d. data (Chen, Li, and Li 2020).
Challenge 2: Local Overfitting and Client Drift. While AdamW accelerates convergence through its adaptivity, it may exacerbate local overfitting. In FL, where each client minimizes its own local objective $f_{i}(\cdot)$, creating a natural gap between the local and global optima. Adaptive optimizers such as AdamW, with stronger update magnitudes, can drive clients further toward their local optima—diverging from the global direction. This leads to client drift as illustrated in Figure 2 (b), which manifests as inconsistencies in local models that degrade the global aggregation performance.
@@
We observe that the Hessian matrix in neural networks exhibits an approximate block-diagonal structure with several dense sub-blocks (Collobert 2004; Zhang et al. 2024b) as shown in Figure 3. In such a structure, a single learning rate can effectively capture the curvature within each block. Leveraging this, we propose a communication-efficient strategy that partitions the second-moment estimate $\boldsymbol{v}$ into $B$ blocks and transmits only the mean of each in Figure 4:
@@
We group the parameters into semantically aligned classes that exhibit similar curvature patterns:
  • Class 1: query and key. Query and Key parameters. Each block corresponds to one attention head, as shown in Figure 3.
@@
These assumptions are standard in federated adaptive optimization literature (Fan et al. 2024; Sun et al. 2023).
@@
The proof is provided in Appendix B. Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$, where $n$ is the number of tatal data, consistent with classical results from PAC theory, stability, and uniform convergence (Shalev-Shwartz and Ben-David 2014). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b(\sigma_{i}^{1/2}+\lambda)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}(\sigma_{i}^{1/2}+\lambda)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error, as shown in Table Table 6 below. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$).
@@
Datasets. We evaluate FedAdamW on both vision and language tasks. (i) For image classification, we use CIFAR-100 (Krizhevsky, Hinton et al. 2009), and Tiny ImageNet (Le and Yang 2015). (ii) For NLP tasks, we adopt benchmark datasets from the GLUE benchmark, including SST-2 (Socher et al. 2013), QQP (Socher et al. 2013). To simulate data heterogeneity across clients, we follow the Dirichlet partitioning scheme (Hsu, Qi, and Brown 2019), where a Dir-0.6 corresponds to a low heterogeneity and Dir-0.1 implies high heterogeneity.
Model Architectures. We explore a variety of model types: (i) ResNet-18 (He et al. 2016) as a representative convolutional neural network (CNN), (ii) Swin Transformer (Liu et al. 2021) and ViT-Tiny (Dosovitskiy et al. 2020) for Vision Transformers, and (iii) RoBERTa-Base (Liu et al. 2019) for large-scale language model.
Baselines. We compare our method against state-of-the-art FL algorithms: FedAvg (McMahan et al. 2017), SCAFFOLD (Karimireddy et al. 2020), FedCM (Xu et al. 2021), FedAdam (Reddi et al. 2020), FedLADA (Sun et al. 2023), Local Adam and Local AdamW.
@@
Questions. Our experiments are designed to answer the following: Q1. Does Local AdamW outperform Local SGD when training Transformer models? Q2. Can FedAdamW effectively address the three challenges identified for AdamW in FL? Q3. Is FedAdamW generally effective across both CNNs and Transformers? Q4. Are individual components of FedAdamW—such as global update correction, decoupled weight decay, and block-wise $\boldsymbol{v}$ averaging—empirically beneficial? Q5. Do our theoretical findings (Theorems 1 and 2) align with empirical results?
@@
Table 4: Ablation study of FedAdamW on CIFAR-100 using ViT-Tiny (Dir-0.1, 300 rounds).
@@
Table 5: Impact of $\alpha$ on FedAdamW using ViT-Tiny on CIFAR-100 (Dir-0.1).
@@ Conclusion
In this work, we proposed a novel federated optimization algorithm (FedAdamW) for training large-scale Transformer models. FedAdamW tackles the key challenges of applying AdamW in federated settings, including high variance in second-moment estimates, local overfitting under non-i.i.d. data, and inefficiencies from frequent reinitialization. It integrates second-moment aggregation, global update correction, and decoupled weight decay. We provided convergence analysis under non-convex and use the PAC Bayesian theory to support its generalization benefits. Extensive experiments on vision and language tasks verified that FedAdamW consistently outperforms strong FL baselines, especially on Transformer architectures, demonstrating its practical and theoretical strengths. We believe FedAdamW opens a new direction for adapting modern optimizers to FL such as LAMB (Chen et al. 2023) or Lion (Chen et al. 2023).
@@ Appendix B Appendix B: Generalization Analysis
Analysis on hypothesis posterior. According to the classical PAC-Bayesian theory (McAllester 1998), the generalization error bound is closely related to the hypothesis posterior learned by an algorithm. We therefore begin by analyzing the hypothesis posterior induced by FedAdamW, followed by its impact on generalization. Following prior work (Xie et al. 2021; Mandt, Hoffman, and Blei 2016; Chaudhari and Soatto 2018; Jastrzbski et al. 2017; Zhu et al. 2018; Zhou et al. 2020, 2024), we adopt the stochastic differential equation (SDE) view of optimization, as it captures similar convergence behavior to that of the actual algorithm. We focus on a simplified version of the second-moment estimate used in AdaGrad (Duchi, Hazan, and Singer 2011), which is given by: $\mathrm{diag}\left(\boldsymbol{n}_{r+1}\right)$, which is commonly used in practice. The observed fisher information matrix, $\frac{1}{n}\sum_{i=1}^{n}\nabla F\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)\nabla F\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)^{\top}$ serves as a reliable approximation of the local Hessian matrix $\boldsymbol{H}_{\boldsymbol{x}^{r}}$ (Pawitan 2001; Jastrzbski et al. 2017; Zhu et al. 2018). Based on these insights, we follow previous works and treat $\boldsymbol{n}_{r}$ as a proxy for the Hessian matrix in our analysis.
@@
Let $\boldsymbol{u}^{r}=\nabla F(\boldsymbol{x}^{r})-\frac{1}{N}\sum_{i=1}^{N}\nabla F_{i}(\boldsymbol{x}^{r})$ denote the heterogeneity gradient noise, and define $\boldsymbol{Q}_{r}=\boldsymbol{n}^{-1/2}$ for simplicity. In the FedAdamW formulation, we omit the small constant $\epsilon$ for clarity, as it does not affect subsequent results. Following (Mandt, Hoffman, and Blei 2017; Smith and Le 2017; Chaudhari and Soatto 2018; Zhu et al. 2018), we assume the noise $\boldsymbol{u}^{r}$ follows a Gaussian distribution $\mathcal{N}(\mathbf{0},\boldsymbol{C}_{\boldsymbol{x}^{r}})$ by the Central Limit Theorem. Under this assumption, the SDE of FedAdamW can be expressed as:
@@
where $N$ is the number of all clients, and $S$ is participate clients. Since we analyze the local convergence around an optimum, this means $\nabla F\left(\boldsymbol{x}^{r}\right)\approx 0$ and the variance of the gradient noise would dominate (Mandt, Hoffman, and Blei 2017; Smith and Le 2017; Chaudhari and Soatto 2018) . So following these works, we approximate $\boldsymbol{C}_{\boldsymbol{x}^{r}}$ as
@@
Building on the posterior analysis, we adopt the PAC-Bayesian framework (McAllester 1998) to study the generalization performance of FedAdamW. Given an algorithm $\mathcal{A}$ and a training set $\mathcal{D}_{\text{tr}}$ with samples $\boldsymbol{\xi}\sim\mathcal{D}$, the learned posterior hypothesis follows a distribution $\mathcal{P}\sim\mathcal{N}(\boldsymbol{x}_{*},\boldsymbol{M}_{\text{{FedAdamW}}})$ as shown in Lemma 7. We define the expected risk as $\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$ and the empirical risk as $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{\mathrm{tr}},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$. In practice, the prior is often chosen as a Gaussian $\mathcal{P}_{\text{pre}}\sim\mathcal{N}(\mathbf{0},\rho\boldsymbol{I})$ (Szegedy et al. 2015; He et al. 2016; Dosovitskiy et al. 2020), due to the lack of prior knowledge about the posterior. Based on Lemma 7, we establish the generalization bound in Theorem 4.
@@
Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$ (up to constant factors), consistent with classical results from PAC theory, stability, and uniform convergence (Vapnik 2006). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b\left(\sigma_{i}^{1/2}+\lambda\right)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}\left(\sigma_{i}^{1/2}+\lambda\right)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error. Since the empirical risk $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{\mathrm{tr}},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$ is typically small for modern over-parameterized models, adjusting $\lambda$ helps reduce the expected risk. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$). To the best of our knowledge, this is the first theoretical result that explicitly demonstrates how weight decay improves the generalization in FL.

### v2

@@ Introduction
With the rapid growth of data and rising concerns over user privacy, traditional centralized training paradigms have become inadequate. Federated Learning (FL) (mcmahan2017communication) offers a scalable and privacy-preserving framework that enables collaborative model training across decentralized clients without sharing raw data (bian2025fedalt; wang2025adaptive; NEURIPS2024_a11e42a3; Bian_2025_ICCV; 10546478). As data becomes increasingly siloed, FL is a practical solution for large-scale distributed deep learning (li2025multi; li2023ultrare; li2024survey; an2022numerical; an2024inspired; an2024robust; liu2023single; liu2025fedadamw; liu2025dp; liu2025fedmuon).
However, recent trends in model design-particularly the rise of large-scale architectures such as GPT (radford2018improving), RoBERTa (liu2019roberta), and Vision Transformers (ViT) (dosovitskiy2020image)—pose new challenges for existing FL algorithms. Specifically, the widely-used FedAvg algorithm, which relies on stochastic gradient descent (SGD(bottou2010large) in local, struggles to efficiently train Transformer models. This is due to the slow convergence and poor adaptivity of SGD in Transformer models (zhang2024transformers; liuimproving), which have more complex architectures compared to CNNs. For example, components such as query, key, and value often require different learning rates to be trained effectively (zhang2024transformers). In contrast, AdamW (loshchilov2017fixing), an adaptive optimizer with decoupled weight decay, has demonstrated superior performance in centralized training of large models based on Transformer (vaswani2017attention; liu2019roberta), offering faster convergence and improved generalization, compared to Adam (kingma2014adam) and SGD.
Empirically, we also observe this advantage in FL: as shown in Figure 1, local training with AdamW (Local AdamW) converges significantly faster than Local SGD  (mcmahan2017communication) for training various Transformer models . However, naively applying AdamW in FL leads to the following new challenges:
@@
Figure 1: Performance of Local SGD and Local AdamW. For training ViT-Base, GPT2, and BERT (liu2019roberta), we carefully tune the learning rate. For training all these Transformer models, Local SGD is still significantly worse than Local AdamW.
@@
  • We propose FedAdamW, a principled FL algorithm tailored for adaptive optimizers. To address the above challenges, FedAdamW integrates global update estimate into local updates to mitigate overfitting and improve consistency. Inspired by the Hessian structure, we design a communication-efficient aggregation strategy that communicates the mean of second-moment across clients.
@@
$\bullet$ Heterogeneity Issues in Federated Learning. Data heterogeneity across clients is a fundamental challenge in FL. A range of algorithms have been proposed to mitigate the adverse effects of non-i.i.d. data distributions. For example, FedProx (li2018federated) introduces a proximal term to restrict local updates; SCAFFOLD (karimireddy2020scaffold) applies control variates to correct client drift; and FedCM (xu2021fedcm) leverages client momentum to stabilize updates. FedNSAM (liu2025consistency) analyzed the consistency between local and global flatness, FedBCGD  (liu2024fedbcgd) proposed a communication-efficient accelerated block coordinate gradient method. FedSWA  (liuimproving) further improved generalization under highly heterogeneous data via stochastic weight averaging.
@@
$\bullet$ Adaptive Optimization in Centralized Settings. Adaptive gradient methods have demonstrated superior empirical performance over SGD in centralized settings, particularly for deep neural networks. Pioneering works include Adagrad (duchi2011adaptive), Adadelta (zeiler2012adadelta), Adam (kingma2014adam), AMSGrad (reddi2019convergence), and AdamW (loshchilov2017fixing). AdamW, in particular, decouples weight decay from gradient updates, offering improved generalization and training stability—attributes especially critical for Transformer models (liu2019roberta; zhang2024transformers; DBLP:journals/tnn/ZhangHLDCW25; zhang2025molebridge; DBLP:conf/aaai/Zhang0LXCCW25; DBLP:conf/aaai/ZhangLXCW024; ding2025traffic; zeng2025FSDrive; zeng2025janusvln; zhang2025fedagc; ENCODER; FineCIR; OFFSET; zhao2024balf; zhao2023benchmark; liao2025convex; liao2024globalpointer; chen2025improving; dang2023ticoserec; dang2025efficient; li2025dual; yang2025distillation; yang2023adaptive; qian2024bird; qian2025vovtrack; limva; li2024savsr; zhou2023fastpillars; pillarhist; zhou2024information; zhou2024lidarptq; ouyang2025multi; wei2025multiple; wei2025comprehensive; ouyang2025graph; ouyang2025dynamic; ouyang2025beaver).
@@
$\bullet$ Adaptive Optimization in Federated Learning. Recent efforts have explored integrating adaptive methods into FL. FedOpt (reddi2020adaptive) incorporates server-side adaptivity using Adam and Yogi. FAFED (wu2023faster) aggregates both the first- and second-moment estimates of Adam across clients to stabilize training. FedAMS  (chen2020toward) shows that averaging the second-moment estimate of Adam is crucial to prevent divergence. More recently, sun2023efficient proposed to only aggregate the second-moment estimate to reduce communication overhead. However, these works only conducted experiments on CNN models. These studies are all based on Adam, which performs poorly with large weight decay.
FL Problem Setup
Challenges of AdamW in FL
Despite the widespread use of AdamW (loshchilov2017fixing; vaswani2017attention) in centralized deep learning, its adaptation to federated settings remains largely unexplored. In this section, we analyze three fundamental challenges that hinder its effectiveness in FL settings.
where $\boldsymbol{v}^{r,k}_{i}$ denotes the second-moment estimate maintained by client $i$ at local step $k$ of round $r$, $\boldsymbol{g}^{r,k}_{i}$ is the stochastic gradient, $\beta_{2}=0.999$ is the exponential decay rate for the second moment, and $\odot$ represents the element-wise (Hadamard) product in Algorithm 1. In FL, data heterogeneity leads to gradient heterogeneity. The squared stochastic gradients $\boldsymbol{g}^{r,k}_{i}\odot\boldsymbol{g}^{r,k}_{i}$ in Eq. (2) amplify the variance of $\boldsymbol{v}$ across clients in Figure 2 (a). This can cause instability and inefficient aggregation, especially when using non-i.i.d. data (chen2020toward).
Challenge 2: Local Overfitting and Client Drift. While AdamW accelerates convergence through its adaptivity, it may exacerbate local overfitting. In FL, where each client minimizes its own local objective $f_{i}(\cdot)$, creating a natural gap between the local and global optima. Adaptive optimizers such as AdamW, with stronger update magnitudes, can drive clients further toward their local optima—diverging from the global direction. This leads to client drift as illustrated in Figure 2 (b), which manifests as inconsistencies in local models that degrade the global performance.
@@
We observe that the Hessian matrix in neural networks exhibits an approximate block-diagonal structure with several dense sub-blocks (collobert2004large; zhang2024adam) as shown in Figure 3. In such a structure, a single learning rate can effectively capture the curvature within each block. Leveraging this, we propose a communication-efficient strategy that partitions the second-moment estimate $\boldsymbol{v}$ into $B$ blocks and transmits only the mean of each in Figure 4:
@@
We group the parameters into semantically aligned classes that exhibit similar curvature patterns as shown in Figure 3:
  • Class 1: query and key. Query and Key parameters. Each block corresponds to one attention head.
@@
These assumptions are standard in federated adaptive optimization literature (fan2024locally; sun2023efficient).
@@
The proof is provided in Appendix B. Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$, where $n$ is the number of tatal data, consistent with classical results from PAC theory, stability, and uniform convergence (shalev2014understanding). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b(\sigma_{i}^{1/2}+\lambda)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}(\sigma_{i}^{1/2}+\lambda)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error, as shown in Table Table 6 below. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$).
@@
Datasets. We evaluate FedAdamW on both vision and language tasks. (i) For image classification, we use CIFAR-100 (krizhevsky2009learning), and Tiny ImageNet (le2015tiny). (ii) For NLP tasks, we adopt benchmark datasets from the GLUE benchmark, including SST-2 (socher2013recursive), QQP (socher2013recursive). To simulate data heterogeneity across clients, we follow the Dirichlet partitioning scheme (hsu2019measuring), where a Dir-0.6 corresponds to a low heterogeneity and Dir-0.1 implies high heterogeneity.
Model Architectures. We explore a variety of model types: (i) ResNet-18 (he2016deep) as a representative convolutional neural network (CNN), (ii) Swin Transformer (liu2021swin) and ViT-Tiny (dosovitskiy2020image) for Vision Transformers, and (iii) RoBERTa-Base (liu2019roberta) for large-scale language model.
Baselines. We compare our method against state-of-the-art FL algorithms: FedAvg (mcmahan2017communication), SCAFFOLD (karimireddy2020scaffold), FedCM (xu2021fedcm), FedAdam (reddi2020adaptive), FedLADA (sun2023efficient), Local Adam and Local AdamW.
@@
textbfQuestions. Our experiments are designed to answer the following: Q1. Does Local AdamW outperform Local SGD when training Transformer models? Q2. Can FedAdamW effectively address the three challenges identified for AdamW in FL? Q3. Is FedAdamW generally effective across both CNNs and Transformers? Q4. Are individual components of FedAdamW—such as global update correction, decoupled weight decay, and block-wise $\boldsymbol{v}$ averaging—empirically beneficial? Q5. Do our theoretical findings (Theorems 1 and 2) align with empirical results?
@@
Table 4: Ablation study of on CIFAR-100 using ViT-Tiny (Dir-0.1, 300 rounds).
@@
Table 5: Impact of $\alpha$ using ViT-Tiny on CIFAR-100 (Dir-0.1).
@@ Conclusion
In this work, we proposed a novel federated optimization algorithm (FedAdamW) for training large-scale Transformer models. FedAdamW tackles the key challenges of applying AdamW in federated settings, including high variance in second-moment estimates, local overfitting under non-i.i.d. data, and inefficiencies from frequent reinitialization. It integrates second-moment aggregation, global update correction, and decoupled weight decay. We provided convergence analysis under non-convex and use the PAC Bayesian theory to support its generalization benefits. Extensive experiments on vision and language tasks verified that FedAdamW consistently outperforms strong FL baselines, especially on Transformer architectures, demonstrating its practical and theoretical strengths. We believe FedAdamW opens a new direction for adapting modern optimizers to FL such as LAMB (chen2023symbolic) or Lion (chen2023symbolic).
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62276182), Peng Cheng Lab Program (No. PCL2023A08), Tianjin Natural Science Foundation (Nos. 24JCYBJC01230, 24JCYBJC01460), and Tianjin Municipal Education Commission Research Plan (No. 2024ZX008).
@@ Appendix B Appendix B: Generalization Analysis
Analysis on hypothesis posterior. According to the classical PAC-Bayesian theory (mcallester1998some), the generalization error bound is closely related to the hypothesis posterior learned by an algorithm. We therefore begin by analyzing the hypothesis posterior induced by FedAdamW, followed by its impact on generalization. Following prior work (xie2021positive; mandt2016variational; chaudhari2018stochastic; jastrzkebski2017three; zhu2018anisotropic; zhou2020towards; zhou2024towards), we adopt the stochastic differential equation (SDE) view of optimization, as it captures similar convergence behavior to that of the actual algorithm. We focus on a simplified version of the second-moment estimate used in AdaGrad (duchi2011adaptive), which is given by: $\mathrm{diag}\left(\boldsymbol{n}_{r+1}\right)$, which is commonly used in practice. The observed fisher information matrix, $\frac{1}{n}\sum_{i=1}^{n}\nabla F\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)\nabla F\left(\boldsymbol{x}^{r};\boldsymbol{\xi}_{i}\right)^{\top}$ serves as a reliable approximation of the local Hessian matrix $\boldsymbol{H}_{\boldsymbol{x}^{r}}$ (pawitan2001all; jastrzkebski2017three; zhu2018anisotropic). Based on these insights, we follow previous works and treat $\boldsymbol{n}_{r}$ as a proxy for the Hessian matrix in our analysis.
@@
Let $\boldsymbol{u}^{r}=\nabla F(\boldsymbol{x}^{r})-\frac{1}{N}\sum_{i=1}^{N}\nabla F_{i}(\boldsymbol{x}^{r})$ denote the heterogeneity gradient noise, and define $\boldsymbol{Q}_{r}=\boldsymbol{n}^{-1/2}$ for simplicity. In the FedAdamW formulation, we omit the small constant $\epsilon$ for clarity, as it does not affect subsequent results. Following (mandt2017stochastic; smith2017bayesian; chaudhari2018stochastic; zhu2018anisotropic), we assume the noise $\boldsymbol{u}^{r}$ follows a Gaussian distribution $\mathcal{N}(\mathbf{0},\boldsymbol{C}_{\boldsymbol{x}^{r}})$ by the Central Limit Theorem. Under this assumption, the SDE of FedAdamW can be expressed as:
@@
where $N$ is the number of all clients, and $S$ is participate clients. Since we analyze the local convergence around an optimum, this means $\nabla F\left(\boldsymbol{x}^{r}\right)\approx 0$ and the variance of the gradient noise would dominate (mandt2017stochastic; smith2017bayesian; chaudhari2018stochastic) . So following these works, we approximate $\boldsymbol{C}_{\boldsymbol{x}^{r}}$ as
@@
Building on the posterior analysis, we adopt the PAC-Bayesian framework (mcallester1998some) to study the generalization performance of FedAdamW. Given an algorithm $\mathcal{A}$ and a training set $\mathcal{D}_{\text{tr}}$ with samples $\boldsymbol{\xi}\sim\mathcal{D}$, the learned posterior hypothesis follows a distribution $\mathcal{P}\sim\mathcal{N}(\boldsymbol{x}_{*},\boldsymbol{M}_{\text{{FedAdamW}}})$ as shown in Lemma 7. We define the expected risk as $\mathbb{E}_{\boldsymbol{\xi}\sim\mathcal{D},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$ and the empirical risk as $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{\mathrm{tr}},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$. In practice, the prior is often chosen as a Gaussian $\mathcal{P}_{\text{pre}}\sim\mathcal{N}(\mathbf{0},\rho\boldsymbol{I})$ (szegedy2015going; he2016deep; dosovitskiy2020image), due to the lack of prior knowledge about the posterior. Based on Lemma 7, we establish the generalization bound in Theorem 4.
@@
Theorem 2 shows that the generalization error of FedAdamW can be upper bounded by $\mathcal{O}(1/\sqrt{n})$ (up to constant factors), consistent with classical results from PAC theory, stability, and uniform convergence (vapnik2006estimation). We further analyze the impact of the decoupled weight decay parameter $\lambda$ on this bound. As $\lambda$ increases, the first term $\sum_{i=1}^{d}\log 2\rho b\left(\sigma_{i}^{1/2}+\lambda\right)\eta^{-1}$ increases, while the second term $\frac{\eta}{2\rho b}\sum_{i=1}^{d}\left(\sigma_{i}^{1/2}+\lambda\right)^{-1}$ decreases. Although choosing the optimal $\lambda$ is challenging in practice, this trade-off suggests that tuning $\lambda$ appropriately can lead to a smaller generalization error. Since the empirical risk $\mathbb{E}_{\boldsymbol{\xi}\in\mathcal{D}_{\mathrm{tr}},\boldsymbol{x}\sim\mathcal{P}}[f(\boldsymbol{x},\boldsymbol{\xi})]$ is typically small for modern over-parameterized models, adjusting $\lambda$ helps reduce the expected risk. This explains why FedAdamW often outperforms Local Adam (which corresponds to $\lambda=0$). To the best of our knowledge, this is the first theoretical result that explicitly demonstrates how weight decay improves the generalization in FL.